OpenStudy (anonymous):

what is an equation of the circle whose diameter ab has endpoints a (-4,2) and b(4,4)

OpenStudy (amistre64):

stack your points subtract them and you get your distances on x and y for the pythag

OpenStudy (amistre64):

-4,2 -(4,4) ------ -8, -2; x = 8 and y = 2 distance = sqrt(x^2 +y^2)

OpenStudy (amistre64):

sqrt(64 +4) = sqrt(68)

OpenStudy (anonymous):

or find the midpoint. midpoint is \[(\frac{-4+4}{2},\frac{2+4}{2})\] i.e. (0,3)

OpenStudy (amistre64):

pull the 4 out and get 2sqrt(17) divide in half to get radius

OpenStudy (amistre64):

rad = sqrt(17)

OpenStudy (amistre64):

middies good for center point of equation yes ;)

OpenStudy (amistre64):

(x-Cx)^2 +(y-Cy)^2 = 17 then

OpenStudy (anonymous):

now find the distance from the midpoint to one endpoint. forget about taking the square root because you formula will have a square in it. \[(4-0)^2+(4-3)^2=16+1=17\]

OpenStudy (anonymous):

so your equation is \[(x-0)^2+(y-3)^2=17\] better known as \[x^2+(y-3)^2=17\]

OpenStudy (anonymous):

hello amistre. been here for a while?

OpenStudy (anonymous):

THANXZ

OpenStudy (amistre64):

i been here since about well..... lost track

OpenStudy (anonymous):

-4,2 -(4,4) ------ -8, -2; x = 8 and y = 2 distance = sqrt(x^2 +y^2)